OpenStudy (anonymous):

Find the equation of the circle passing through the points (-4,-2) and (2,0) and having its center on the line 5x-2y=19. I need two forms of answer... general and standard... its okay if you give me a general equation b'coz I knew already how to convert general equation to standard.

OpenStudy (anonymous):

how is your work finding last digit of pi ?

OpenStudy (anonymous):

wow... until now you still remembered it?

OpenStudy (anonymous):

wow indeed lol

OpenStudy (angela210793):

the eq of the circle will b (x-x1)^2+(y-y1)^2=r^2 since the points given r in the circle they prove the eq of the circle and since the center is in the line 5x-2y=19 then its coordinates prove the eq of the circle... so we can form the system (-4-x1)^2+(-2-x1)=r^2 (2-x1)^2+(0-y1)^2=r^2 5x1-2y1=19

OpenStudy (anonymous):

what about the general eq.

OpenStudy (angela210793):

hmm....isn't this the general eq?

OpenStudy (anonymous):

the general eq. is this x^2+y^2+Dx+Ey+F=0

OpenStudy (angela210793):

well after u've found the eq (x-x1)^2+(y-y1)^2=r^2 all u have to do is transform it x^2-2xx1-x1^2...etc...and this is how i get it

OpenStudy (anonymous):

the standard eq. is this\[(x+h)^{2}+(y+k)^{2}=0\]

OpenStudy (anonymous):

I think you've used the calculus in solving this problem.

OpenStudy (anonymous):

I think that's conic

OpenStudy (angela210793):

tht's conic....

OpenStudy (anonymous):

That's a sort of "help" formula based on completing the square..

OpenStudy (anonymous):

oh my brain malfunctioned that is the general equation sorry

OpenStudy (anonymous):

Your brain didn't malfunction, it isn't standard, just meant to help the student...

OpenStudy (anonymous):

Thanks estudier : )

OpenStudy (anonymous):

I don't like questions like this ...

OpenStudy (anonymous):

I still think the simplest method for solving this sort of thing is to translate to the origin but I suppose that might be going beyond....

OpenStudy (anonymous):

eaularch just use the formulas distance one then convert it into any equation you like

OpenStudy (anonymous):

what is that is that really way beyond translate

OpenStudy (anonymous):

okay you're saying make the center 0,0 right?

OpenStudy (anonymous):

Everything is just much easier to calculate at (0,0)..

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Then when u got it, translate it back..

OpenStudy (anonymous):

Anyway, I am just trying to think of a simple way to explain...(without translation).

OpenStudy (anonymous):

yeah thanks I will try this tonight thanks estudier these type of questions are asked a lot in india especially in IIT,AIEEE exams (for entering top colleges in India) I will try this. Thanks I gotta go now bye and thanks estudier

OpenStudy (anonymous):

Ciao!

OpenStudy (angela210793):

Estudier it is easier if the center was at (0;0) but it is not ;/

OpenStudy (anonymous):

It's just a translation, quite simple.